The spot price per share is $115 and the risk free rate is 5% per annum on a continuously compounded basis. The annual volatility is 20% and the stock does not pay any dividend. All options have a one-year maturity. In answering the questions below use a binomial tree with three steps. Each step should be one-third of a year.

1)Using the binomial tree, compute the price at time 0 of a one-year European put option on 100 shares of stock with a strike price of $115 per share.
Answer #1

For a Binomial Model, The value of a Put option under the risk neutral probability measure is defined as the Discounted value of the expected payoff.

Or

Put option value Pt= e-rt * EQ[Payoff(XT)]

BINOMIAL PRICING (“TREE) MODEL Risk-neutral probabilities erat – d Up-step probability => , u-d where u zpovar+qar and dze-o

Here, u= Up step factor

d= Down step factor

q= up step probability(as defined above)

The following Put price has been computed on excel:

sd 20% dt= 0.33
Rf 5% u 1.122401 =EXP(B23*SQRT(E23)) Share Price
Spot Price(St) 115.00 d 0.890947 =1/E24 Payoff
Strike(K) 115 q 0.543777 =(EXP(B24*E23)-E25)/(E24-E25)
t=0 t=1 t=2 t=3
Formulas
162.608 =F35*$E$24
0 =MAX($B$26-G33,0)
144.8751 =E37*$E$24
0 =EXP(-$B$24*$E$23)*(G34*$E$26+G38*(1-$E$26))
129.0761 129.0761 =F39*$E$24
2.524719 0 =MAX($B$26-G37,0)
115.00 115 =E41*$E$24
Pt= 7.091836 5.626959 =EXP(-$B$24*$E$23)*(G38*$E$26+G42*(1-$E$26))
102.4589 102.4589 =F43*$E$24
12.79667 12.54107 =MAX($B$26-G41,0)
91.28551 =E41*$E$25
21.81371 =EXP(-$B$24*$E$23)*(G42*$E$26+G46*(1-$E$26))
81.33057 =F43*$E$25
33.66943 =MAX($B$26-G45,0)

The price for put on 1 share is Pt= 7.091836.

Therefore price for Put on 100 shares is Pt=709.1836.